OpenStudy (anonymous):

int x^2/sqrt(9-x^2) please help :)

OpenStudy (jhannybean):

Seems like a trig sub.

OpenStudy (anonymous):

i got x=3sintheta dx=3costhetadtheta

OpenStudy (anonymous):

so \[\int\limits_{}^{}9\sin^2\theta/\sqrt{9-9\sin^2\theta}(3\cos \theta d \theta)\]

OpenStudy (anonymous):

now where im lost is the book says "simplify" and then makes it into \[9\int\limits_{}^{}\sin^2\theta d \theta\]

OpenStudy (anonymous):

with no explanation of what happened to everything else and i don't know how this was simplified?

OpenStudy (aum):

\[ \int \frac{9\sin^2\theta}{\sqrt{9-9\sin^2\theta}}3\cos \theta d \theta = \int \frac{9\sin^2\theta}{\sqrt{9(1-\sin^2\theta)}}3\cos \theta d \theta = \\ \int \frac{9\sin^2\theta}{\sqrt{9\cos^2\theta)}}3\cos \theta d \theta = \int \frac{9\sin^2\theta}{3\cos\theta}3\cos \theta d \theta = \int 9\sin^2(\theta)d\theta \]

OpenStudy (anonymous):

AHH ok thank you so much

OpenStudy (aum):

You are welcome.

OpenStudy (anonymous):

ok just one thing though if you have sqrt 9cos^2theta

OpenStudy (jhannybean):

Ahh he beat me to it! haha. Yeah you were on the right path :D

OpenStudy (anonymous):

ahhh you square rooted it

OpenStudy (anonymous):

you sly devil you

OpenStudy (jhannybean):

You can factor out the 9 and treat cos theta as a constant, therefore pulling out of one them you'll have 3 cos theta.